OpenStudy (anonymous):

help me to answer this : A 75 kg stuntman drops from a tree onto a 50 kg sled that is moving on a frozen lake with a velocity of 10 m/s towards the shore. how can i find the speed of sled after stuntman is on board?

OpenStudy (chmvijay):

see this is based on momentum problem right :)

OpenStudy (chmvijay):

problem says 50 kg sled that is moving on a frozen lake with a velocity of 10 m/s mean it has momentum of MV right we call it as Ms*Vs

OpenStudy (chmvijay):

and when a man of 75 Kg falls on it its momentum get disturbed now he will have a new momentum that is 75+50 and Velocity we don't know we can now have new momentum that is 75+50 8 Vf

OpenStudy (chmvijay):

equate both for the conservation of momentum

OpenStudy (chmvijay):

Ms Vs = (Ms+Mman)Vf 50*10 =(50+75)Vf 500/125 = 4 = vf

OpenStudy (chmvijay):

speed is 4 m/s

OpenStudy (anonymous):

ouh..i think this question need to use frmula 1/2 mv2..because tit has mass and velocity

OpenStudy (chmvijay):

noo i dont think soo its based on momentum and not kientic energy :)

OpenStudy (anonymous):

yaa...i got it...thanks for helping..

OpenStudy (chmvijay):

ur welcome :)